OpenStudy (anonymous):

Evaluate the line integral (9x+7y^3) ds, where the curve C is r(t) = (t-9)i + tj, 0<=t<=1. I'm thinking I input the boundaries of t into r(t) and the put them into the integral...? I will follow and give medal!

OpenStudy (anonymous):

The integral looks like this: \[\int\limits_{C}^{}(9x+7y^3)ds\]

OpenStudy (anonymous):

I agree

OpenStudy (irishboy123):

you can do this in a number of ways. if you note that \(r(t) = <t-9, t>\), then you can see that \(x = t-9, y = t\) so \(y = x+9\) that takes you straight into Cartesian, so you can use \(ds = \sqrt{y'^1 + 1 } \; dx\)

OpenStudy (irishboy123):

or you can follow the paramaterisation route. maybe that is what you are learning as it seems to be quite an endemic ... try both :-)

OpenStudy (anonymous):

Oh yeah. His response is better cause when you break it down, you get your answer like I am doing in my notebook right now

OpenStudy (anonymous):

I've got it!!!

OpenStudy (anonymous):

So ds = \[\sqrt{y+1}\]?

OpenStudy (anonymous):

I know the answer guys, if you want it

OpenStudy (irishboy123):

sorry , bad typo \(ds = \sqrt{\color{red}{(y')^2} + 1 } \; dx\) and \(y' = 1\) as \(y = x+9\)

OpenStudy (anonymous):

Ah okay, so ds = sqrt(2)

OpenStudy (anonymous):

bye0

OpenStudy (anonymous):

bye

OpenStudy (irishboy123):

or you can paramaterise so you have \(ds = \sqrt{\dot x ^2 + \dot y ^2} \; dt\) you should get the same answer, try it

OpenStudy (anonymous):

Okay, thank you very much, Irish

OpenStudy (anonymous):

You guys don't need help so i'm going to go I was going to explain how i got the answer that I got and break it down for you guys so you can understand a little better but you are ignoring me so see yah later

OpenStudy (anonymous):

Ignorers:p

OpenStudy (irishboy123):

well, army, come back if it crashes. morgan, no offense was intended, seriously. post your solution & ideas, pls.

OpenStudy (anonymous):

I'm going to stay only if you need my help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well, I sorta agree with you so I just realized what you wrote so yeah. Sorry for being rude. I agree with Irish

OpenStudy (anonymous):

That's what I can see to use as a system for the whole thing like you can detect whether the st

OpenStudy (anonymous):

I meant whether the system will need to use a dialect sinatrical order or not

OpenStudy (anonymous):

IDK how to spell sinatrical

OpenStudy (irishboy123):

me neither, lol!! 🍀

OpenStudy (anonymous):

hahaha

OpenStudy (irishboy123):

i don't even know what it means :p

OpenStudy (anonymous):

Hold on a sec... I think it mae

OpenStudy (anonymous):

*means

OpenStudy (anonymous):

It means a polytheistic term of a function in the math category that can differ whether something can be evaluated by a separatism or not. Weird, huh!

OpenStudy (anonymous):

It's awfully confusing